OpenStudy (anonymous):

If f(x)=3x^2-2 and g(x)=4x+2, what is the value of (f+g)(-1) A −7 B −1 C 1 D 7

OpenStudy (danielajohana):

all you have to do here is plug in the equations that you have. (f+g)(-1) (3x^2 -2 + 4x+2)(-1) then you combine like terms (3x^2 + 4x )(-1) then you solve for x and multiply it by -1 and you have your answer.

OpenStudy (anonymous):

Thanks!

OpenStudy (danielajohana):

You're welcome!

OpenStudy (anonymous):

@DanielaJohana I got -3... but I'm not sure what I did wrong... >.<

OpenStudy (danielajohana):

Okay let me try this... give me a moment @potatopeel

OpenStudy (danielajohana):

@potatopeel, I got x=0 and x=-1.3333 (f+g)(-1) f(x)= 3x^2 -2 g(x)= 4x +2 (3x^2-2 + 4x +2)(-1) (3x^2 +4x)(-1) -3x^2 -4x x(-3x-4)=0 x=0 and 0= -3x-4 +4 +4 4=-3x 4/-3=x -1.33333=x